The probabilities that the costumer will buy 1,2,3,4,5 items in a grocery store are 1/6,1/6,1/6,2,/6,1/6 respectively. what is the average number of item that



a costumer will buy?

1
Expert's answer
2022-03-21T11:23:06-0400

"EX=1\cdot \frac{1}{6}+2\cdot \frac{1}{6}+3\cdot \frac{1}{6}+4\cdot \frac{2}{6}+5\cdot \frac{1}{6}=3.16667"